OpenStudy (mew55):

integral 0 to 1 dx/x^2 the dx/x^2 is a fraction and i dont know where to start

OpenStudy (anonymous):

Say that 1/x^2 is x^(-2)

OpenStudy (mew55):

but wut i am suppose to do w/ the x^-2 after that? that is wut i am confused on

OpenStudy (turingtest):

\[\int x^ndx=\frac{x^{n+1}}{n+1}+C\]

OpenStudy (anonymous):

answer is (x^(-2+1))/(-2+1)=-1/x

OpenStudy (mew55):

thatz all there is to it?

OpenStudy (anonymous):

yes